solve system using addition method
x+2y=32
9x-2y=28
solution (?,?)

Answer by rapaljer(4671)

x+2y=32
9x-2y=28

Add the equations together:
10x = 60
x=6

Substitute x=6 into the first equation:
x+2y=32
6+2y=32
2y=32-6
2y=26
y=13

Check in the second equation:
9x-2y=28
9(6)-2(13)=28
54 - 26=28
It checks!!

R^2